# A Comparative Analysis of Wealth Index Predictions in Africa between three Multi-Source Inference Modelsthanks: Please cite the ECMLPKDD’24 version of this paper

### v1

@@ Abstract
Poverty map inference is a critical area of research, with growing interest in both traditional and modern techniques, ranging from regression models to convolutional neural networks applied to tabular data, images, and networks. Despite extensive focus on the validation of training phases, the scrutiny of final predictions remains limited. Here, we compare the Relative Wealth Index (RWI) inferred by Chi et al. (2021) with the International Wealth Index (IWI) inferred by Lee and Braithwaite (2022) and Espín-Noboa et al. (2023) across six Sub-Saharan African countries. Our analysis focuses on identifying trends and discrepancies in wealth predictions over time. Our results show that the predictions by Chi et al. and Espín-Noboa et al. align with general GDP trends, with differences expected due to the distinct time-frames of the training sets. However, predictions by Lee and Braithwaite diverge significantly, indicating potential issues with the validity of the model. These discrepancies highlight the need for policymakers and stakeholders in Africa to rigorously audit models that predict wealth, especially those used for decision-making on the ground. These and other techniques require continuous verification and refinement to enhance their reliability and ensure that poverty alleviation strategies are well-founded.
@@ 1 Introduction
Another critical issue is the choice of ground-truth indicators. Various studies use socioeconomic indicators such as income per capita [7], housing prices [26], and wealth indexes [2, 16, 8], each with its strengths and limitations [20]. Different wealth indexes such as the Relative Wealth Index (RWI), the Comparative Wealth Index (CWI), the Harmonized Wealth Index (HWI), and the International Wealth Index (IWI), provide distinct perspectives on wealth but also come with inherent constraints [16]. This variability makes it challenging to determine the most reliable inferred poverty map when comparing methodologies with similar predictive power.
If the validations of these different models are equally accurate, their resulting poverty maps should theoretically be similar. However, any discrepancies would prompt an investigation into the reasons for these differences and an assessment of which model is the most accurate. In this study, we aim to address this gap by comparing three recent wealth inference approaches: RWI by Chi et al. [2], and IWI by Lee and Braithwaite [16], and by Espín-Noboa et al. [8]. We show their inferred poverty maps across six African countries and examine the degree of concordance of these predictions with established GDP trends, aiming to evaluate the robustness and reliability of these methodologies. The analysis is two-fold. First, we compare the predicted distributions of wealth to assess the overall poverty inferred for each country by each model. Second, considering the differences in geo-located places each approach uses in the final poverty map, we verify whether overlapping locations across methods receive consistent predictions and align with expected poverty trends. In adherence to open science practices, we make both the code and the predictions openly available [9].
@@ 2 Preliminaries
In this study, we compare the predictions of high-resolution poverty maps across six African countries from three different studies: M1 by Chi et al. (2021) [2], M2 by Lee and Braithwaite (2022) [16], and M3 by Espín-Noboa et al. (2023) [8].
@@ 2.1 Ground-Truth Data
Ground-truth data in M1, M2, and M3 are derived from nationally-representative, population-based household surveys conducted by the Demographic and Health Survey (DHS) Program [12]. These surveys encompass asset/wealth information and sub-regional geomarkers. Each household survey includes numerous questions on socioeconomic status. From the responses, DHS calculates the Relative Wealth Index (RWI) [3, 23] using the first principal component derived from a Principal Component Analysis (PCA) [11, 14] of a standardized set of 15 questions concerning assets and housing characteristics. These questions cover various aspects such as electricity, telephone, automobile, refrigerator, TV, water supply, toilet facilities, and construction materials. This RWI serves as the ground-truth measure of wealth in M1. In contrast, M2 and M3 compute the International Wealth Index (IWI) [22], a similar metric, which assesses household material well-being and economic status across low and middle-income countries. The IWI is derived from 12 housing characteristics; M2 excludes two of them, while M3 uses them all. PCA is used to determine asset weights, which are re-scaled to produce IWI scores ranging from 0 to 100, representing the spectrum from no assets to the highest wealth. The IWI’s advantage lies in its comparability across different countries and years, unlike the country-specific RWI used in M1.
The studies also differ in their spatial resolution. M1 employs a 2.4 km grid system defined by Bing tile coordinates to map DHS clusters to grid cells. Conversely, M2 and M3 use actual DHS cluster locations which typically contain 25-30 households, and are geo-located with added noise to protect respondents privacy. The noise varies: up to 2 km for urban and 5 km for rural clusters, with 1% of rural clusters displaced by up to 10 km [1].
Table 1: Years of collected DHS ground-truth and features per study. Instances where two or more models used the same ground-truth data are few: M2 and M3 in Uganda, M1 and M2 in Rwanda, and all models in South Africa. M2 inferred Gabon’s poverty map using a model trained on data from 25 Sub-Saharan African countries, as the most recent survey from 2012 was considered outdated. The table also categorizes the features utilized by each model. Note that for this analysis, as compared to the original models reported in [8], the models of M3 were retrained without features from mobility and demographics, as the mobility data was discontinued and the Facebook Marketing API redefined its reach estimates. This change did not affect performance.
Country M1 M2 M3
Sierra Leone 2013 2016 2016,2019
Liberia 2013 2016 2019,2022
Uganda 2016 2016,2018-19 2016,2018-19
Rwanda 2014-15 2014-15 2019
South Africa 2016 2016 2016
Gabon 2012 Multiple 2019
Feature M1 M2 M3*
@@
Mobility network $\dagger$ - - (✓)
Demographics of people $\dagger$ - - (✓)
@@
Number of total features $\ddagger$ 12+100 249+4 173+784
@@
$\ddagger$ Metadata (non-imagery) features (12, 249, and 173, resp.) plus image-based features (100, 4, and 784, resp.).
The final poverty map is defined by the grid-cells covering the entire country (excluding those used in training) in M1, and by the populated places (i.e., cities, towns, villages, hamlets and isolated dwellings) extracted from OpenStreetMap (OSM) in M2 and M3 (M3 additionally includes neighborhoods).
@@
Figure 1: Available GPD per capita and Gini index. Top: GDP values are aligned with the years of the surveys used in the models (x-axis). Bottom: The most recent available Gini indices are shown for reference. Uganda is the only country with GDP and Gini index data available for approximately the same years as the surveys. Both metric values are provided by the World Bank.
@@ 2.2 Models
M1: RWI inference model by Chi et al. 2021 [2, 5] utilizes $112$ features in total, including high-resolution satellite imagery, mobile phone network data, topographic maps, and aggregated, de-identified connectivity data from Facebook. These diverse data sources are processed through deep learning and computational algorithms to generate quantitative features for each village. These features are then used to train a gradient-boosted regression tree, a supervised machine-learning model, to predict RWI scores. Hyperparameters for the gradient-boosted tree are tuned to minimize the cross-validated mean-squared error (MSE), using a grid search. They use three cross-validation approaches: 5-fold cross-validation (basic CV), leave-one-country-out cross-validation (leave-country-out), and spatially stratified cross-validation (spatial CV). The use of different cross-validation methods helped highlight the potential upward bias in performance due to spatial auto-correlation in the training and test data. All analysis in this study is based on spatial CV, ensuring conservative and appropriate model training for geographic data with spatial auto-correlation. For our comparison analysis, we linearly transform the RWI scores to their IWI equivalents in M3 (see Equation 1) as both are highly correlated (i.e., Pearson’s correlation coefficients: $0.93-0.99$ [16]).
@@
M2: IWI inference model by Lee and Braithwaite 2022 [16] combines feature-based and image-based models to enhance each other’s performance. It uses two cross-validation approaches similar to M1: 5-fold cross-validation (basic CV) and leave-one-country-out cross-validation (leave-country-out). Initially, an XGBoost (eXtreme Gradient Boosting) algorithm is trained on $249$ features (infrastructure, night-time luminosity, and population density) to predict IWI scores. Bayesian optimization is employed to tune the hyperparameters, selecting the best configuration for the training dataset. The XGBoost model’s predictions are then used to train a Convolutional Neural Network (CNN) model that outputs a probability distribution of socioeconomic status (rich, upper-middle class, lower-middle class, poor) for a 1-square-mile area. These probabilities are then fed back into the XGBoost model, creating a cycle where the models iteratively improve by providing better training data.
@@
M3: IWI inference model by Espín-Noboa et al. 2023 [8] extracts $957$ features from seven data sources. Each location is defined by the centroid of each place, using a default bounding-box width of 1 mile. Some features are also queried within 2, 5, and 10 Km to capture the original locations of the DHS clusters, similar to M2. Features include high-resolution satellite imagery, demographics, mobility networks, and mobile phone antennas. Three machine learning models are proposed to predict both the mean and standard deviation of IWI scores, providing insights into the wealth distribution within clusters. The first model is a CNN that predicts IWI scores from daylight satellite images. This CNN comprises 22 layers, with the final layer using a linear activation function and MSE as the loss function, tuning 4 hyperparameters. The second model is a CatBoost regressor (CB), predicting IWI scores from 173 metadata features and evaluated individually for each data source. This model is tuned using 11 hyperparameters. The third model combines CNN+CB by feeding the third-to-last layer of the CNN into the CB model, adding 784 features. For training, data is partitioned into 80% train and 20% test sets, stratified into 10 equal-width bins of wealth. A 4-fold cross-validation on the train set, stratified by SES (ses CV), is used to tune hyperparameters via Random Search on 200 combinations. This process is repeated three times independently with different random seeds, to control for fluctuations, and mean performance is reported. The models are evaluated on various configurations of training data, including different years of ground-truth data, relocation strategies, data augmentation ($*_{a}$), sample weight strategies ($*_{w}$), and feature sources, ensuring robust and comprehensive performance evaluation. For this analysis, we choose the model with best performance (lowest root-mean-square error RMSE) in each country: CB for Sierra Leone and Liberia, CBw for Uganda and Rwanda, and CNNa+CB for Gabon and South Africa.
@@
In summary, all models use a multi-modal and hybrid approach, combining features from various data sources and two model architectures, see Table 1 (bottom). M1 and M3 use satellite images by pre-training a CNN model and feeding the image embeddings into a gradient boosting algorithm. M2 also uses a hybrid approach but first pre-trains a gradient boosting algorithm to train a CNN, and then creates a feedback loop that enhances the initial model. Only M1 estimates the expected error of each location, though its explanatory power is limited (average $R^{2}=0.09$). All models predict the mean wealth of locations, but only M3 additionally predicts their standard deviation, and proposed two additional models to separate the effects of images and metadata.
@@ 2.3 Expected Trends
Before digging into the comparison of predictions between the three models of interest [2, 16, 8], it is crucial to acknowledge the discrepancies in the ground-truth data used to train them. Table 1 (top) outlines the specific years for which each approach utilized data for each country. This variation in time is significant because wealth and poverty are dynamic concepts that can fluctuate over time. To address this, rather than comparing directly the predicted values, we focus on the trends in predictions. We will achieve this by examining the trends of changes in Gross Domestic Product (GDP) per capita provided by the World Bank [27], as depicted in Figure 1 (top). By analyzing these trends, we can hypothesize that the comparison between RWI and IWI will mirror the trends observed in GDP. This approach allows us to account for potential discrepancies arising from the distinct time-frames of the datasets.
@@
Table 2: Prediction statistics. We show the number $N$ of target locations contained in the inferred poverty map, and the mean, standard deviation, and Gini coefficient of the predicted wealth scores for each model. RWI scores are linearly transformed to their IWI equivalents, see Equation 1.
Country M1 ($\hat{IWI}$) M2 (IWI) M3 (IWI)
$N$ Mean SD Gini $N$ Mean SD Gini $N$ Mean SD Gini
Sierra Leone 8435 29.12 4.92 8.79 13040 18.93 10.34 23.01 9881 21.87 6.09 11.54
Liberia 6189 28.62 3.71 6.72 16525 14.57 8.79 25.93 15597 23.35 3.74 7.75
Uganda 25404 41.35 3.06 3.92 27482 22.27 11.26 25.09 27791 41.83 5.21 5.95
Rwanda 3715 45.81 2.93 3.38 23323 24.40 14.51 28.96 1150 47.79 3.46 3.74
South Africa 39072 48.95 8.59 9.69 22687 61.14 18.10 16.85 3043 52.65 14.11 15.04
Gabon 1110 51.40 5.15 5.32 4835 21.89 8.86 17.93 1135 48.32 6.72 7.38
Table 3: Expected vs. obtained differences in wealth. The direction of change in wealth between M1 and M3 generally aligns with World Bank GDP trends, especially in SLE, LBR, UGA, RWA, and GAB. In contrast, the direction of change in wealth between M2 and M3 is only observed in RWA, with larger than expected differences. Most comparisons are significantly different ($p<0.001$, using a two-sample KS-test), except in SLE for M1 and M3. Highlighted values accurately capture the expected trend’s sign (+/-).
Country M1 ($\hat{IWI}$) vs. M3 ($IWI$) M2 ($IWI$) vs. M3 ($IWI$)
@@
Exp.
IWI
Obt.
IWI
Exp.
Gini
Obt.
Gini
KS
Sierra Leone 0.16 0.14 NA -2.75 0.74 0.01 -0.07 NA 11.47 0.62 ***
Liberia 0.04 0.10 NA -1.02 0.67 *** 0.04 -0.23 NA 18.19 0.83 ***
Uganda -0.04 -0.01 -1.7 -2.02 0.13 *** 0.0 -0.31 0.0 19.15 0.87 ***
Rwanda -0.05 -0.02 NA -0.36 0.34 *** -0.05 -0.32 NA 25.22 0.85 ***
South Africa 0.00 -0.04 0.0 -5.34 0.18 *** 0.0 0.07 0.0 1.81 0.30 ***
Gabon 0.11 0.03 NA -2.06 0.45 *** NA -0.38 NA 10.55 0.95 ***
Expected IWI Difference $=(GDP_{Y_{M*}}-GDP_{Y_{M3}})/(GDP_{Y_{M*}}+GDP_{Y_{M3}})$
Obtained IWI Difference $=(\mu_{IWI_{M*}}-\mu_{IWI_{M3}})/(\mu_{IWI_{M*}}+\mu_{IWI_{M3}})$
Expected Gini Difference $=(Gini_{Y_{M*}}-Gini_{Y_{M3}})$
Obtained Gini Difference $=(Gini_{M*}-Gini_{M3})$

### v3

@@ Abstract
Poverty map inference has become a critical focus of research, utilizing both traditional and modern techniques, ranging from regression models to convolutional neural networks applied to tabular data, satellite imagery, and networks. While much attention has been given to validating models during the training phase, the final predictions have received less scrutiny. In this study, we analyze the International Wealth Index (IWI) predicted by Lee and Braithwaite (2022) and Espín-Noboa et al. (2023), alongside the Relative Wealth Index (RWI) inferred by Chi et al. (2022), across six Sub-Saharan African countries. Our analysis reveals trends and discrepancies in wealth predictions between these models. In particular, significant and unexpected discrepancies between the predictions of Lee and Braithwaite and Espín-Noboa et al., even after accounting for differences in training data. In contrast, the shape of the wealth distributions predicted by Espín-Noboa et al. and Chi et al. are more closely aligned, suggesting similar levels of skewness. These findings raise concerns about the validity of certain models and emphasize the importance of rigorous audits for wealth prediction algorithms used in policy-making. Continuous validation and refinement are essential to ensure the reliability of these models, particularly when they inform poverty alleviation strategies.
@@ 1 Introduction
Another critical issue is the choice of ground-truth indicators. Various studies use socioeconomic indicators such as income per capita [7], housing prices [33], and wealth indexes [2, 17, 8], each with its strengths and limitations [23]. Different wealth indexes such as the Relative Wealth Index (RWI) [24], the Comparative Wealth Index (CWI) [25], the Harmonized Wealth Index (HWI) [28], and the International Wealth Index (IWI) [27], provide distinct perspectives on wealth but also come with inherent constraints [17]. This variability makes it challenging to determine the most reliable inferred poverty map when comparing methodologies with similar predictive power.
If the validations of these different models are equally accurate, their resulting poverty maps should theoretically be similar. However, any discrepancies would prompt an investigation into the reasons for these differences and an assessment of which model is more accurate. In this study, we aim to address this gap by comparing three recent wealth inference approaches: IWI by Lee and Braithwaite [17], and by Espín-Noboa et al. [8], and RWI by Chi et al. [2]. We show their inferred poverty maps across six African countries and examine the degree of concordance of these predictions with established GDP trends, aiming to evaluate the robustness and reliability of these methodologies. The analysis is two-fold. First, we compare the predicted distributions of wealth to assess the overall poverty inferred for each country by each model. Second, considering the differences in geo-located places each approach uses in the final poverty map, we verify whether overlapping locations across methods receive consistent predictions and align with expected poverty trends. In adherence to open science practices, we make both the code and the predictions openly available [9].
@@ 2 Preliminaries
In this study, we compare the predictions of high-resolution poverty maps across six African countries from three different studies: M1 by Lee and Braithwaite (2022) [17], M2 by Espín-Noboa et al. (2023) [8], and M3 by Chi et al. (2022) [2].
@@ 2.1 Ground-Truth Data
Ground-truth data in M1, M2, and M3 are derived from nationally-representative, population-based household surveys conducted by the Demographic and Health Survey (DHS) Program [13]. These surveys encompass asset/wealth information and sub-regional geomarkers. Each household survey includes numerous questions on socioeconomic status. From the responses, DHS calculates the Relative Wealth Index (RWI) [3, 29] using the first principal component derived from a Principal Component Analysis (PCA) [12, 15] of a standardized set of 15 questions concerning assets and housing characteristics. These questions cover various aspects such as electricity, telephone, automobile, refrigerator, TV, water supply, toilet facilities, and construction materials. This RWI serves as the ground-truth measure of wealth in M3. In contrast, M1 and M2 compute the International Wealth Index (IWI) [27], a similar metric, which assesses household material well-being and economic status across low and middle-income countries. The IWI is derived from 12 housing characteristics; M1 excludes two of them, while M2 uses them all. PCA is used to determine asset weights, which are re-scaled to produce IWI scores ranging from 0 to 100, representing the spectrum from no assets to the highest wealth. The IWI’s advantage lies in its comparability across different countries and years, unlike the country-specific RWI used in M3.
The studies also differ in their spatial resolution. M1 and M2 use actual DHS cluster locations which typically contain 25-30 households, and are geo-located with added noise to protect respondents privacy. The noise varies: up to 2 km for urban and 5 km for rural clusters, with 1% of rural clusters displaced by up to 10 km [1]. Conversely, M3 employs a 2.4 km grid system defined by Bing tile coordinates to map DHS clusters to grid cells.
Table 1: Years of collected DHS ground-truth and features per study. Instances where two or more models used the same ground-truth data are few: M1 and M2 in Uganda, M1 and M3 in Rwanda, and all models in South Africa. M1 inferred Gabon’s poverty map using a model trained on data from 25 Sub-Saharan African countries, as the most recent survey from Gabon in 2012 was considered outdated. The table also categorizes the features utilized by each model. Note that for this analysis, unlike the original models reported in [8], the M2 models were retrained from scratch without mobility and demographic features, as the mobility data was discontinued and the Facebook Marketing API redefined its reach estimates. This change did not affect performance.
Country (ISO 3166-1 alpha-3 Code) M1 M2 M3
Sierra Leone (SLE) 2016 2016,2019 2013
Liberia (LBR) 2016 2019,2022 2013
Uganda (UGA) 2016,2018-19 2016,2018-19 2016
Rwanda (RWA) 2014-15 2019 2014-15
South Africa (ZAF) 2016 2016 2016
Gabon (GAB) Multiple 2019 2012
Feature M1 M2* M3
@@
Mobility network $\dagger$ - (✓) -
Demographics of people $\dagger$ - (✓) -
@@
Number of total features $\ddagger$ 249+4 173+784 12+100
@@
$\ddagger$ Metadata (non-imagery) features (249, 173, and 12 resp.) plus image-based features (4, 784, and 100 resp.).
The final poverty map is defined by the populated places (i.e., cities, towns, villages, hamlets and isolated dwellings) extracted from OpenStreetMap (OSM) in M1 and M2 (M2 additionally includes neighborhoods), and by the grid-cells covering the entire country in M3.
@@
Figure 1: Available GDP per capita and Gini index. Top: GDP values are aligned with the years of the surveys used in the models (x-axis). Bottom: The most recent available Gini indices are shown for reference. Uganda is the only country with GDP and Gini index data available for approximately the same years as the surveys. Both metric values are provided by the World Bank [34, 35].
@@ 2.2 Models
M1: IWI inference model by Lee and Braithwaite 2022 [17] combines feature-based and image-based models to enhance each other’s performance. It uses two cross-validation approaches: 5-fold cross-validation (basic CV) and leave-one-country-out cross-validation (leave-country-out). Initially, an eXtreme Gradient Boosting (XGBoost) algorithm is trained on $249$ features (infrastructure, night-time luminosity, and population density) to predict IWI scores. Bayesian optimization is employed to tune the hyperparameters, selecting the best configuration for the training dataset. The XGBoost model’s predictions are then used to train a Convolutional Neural Network (CNN) model that outputs a probability distribution of socioeconomic status (rich, upper-middle class, lower-middle class, poor) for a 1-square-mile area. These probabilities are then fed back into the XGBoost model, creating a cycle where the models iteratively improve by providing better training data.
@@
M2: IWI inference model by Espín-Noboa et al. 2023 [8] extracts $957$ features from seven data sources. Each location is defined by the centroid of each place, using a default bounding-box width of 1 mile. Some features are also queried within 2, 5, and 10 Km to capture the original locations of the DHS clusters, similar to M1. Features include high-resolution satellite imagery, demographics, mobility networks, and mobile phone antennas. Three machine learning models are proposed to predict both the mean and standard deviation of IWI scores, providing insights into the wealth distribution within clusters. The first model is a CNN that predicts IWI scores from daylight satellite images. This CNN comprises 22 layers, with the final layer using a linear activation function and mean-squared error (MSE) as the loss function, tuning 4 hyperparameters. The second model is a CatBoost regressor (CB), predicting IWI scores from 173 metadata features and evaluated individually for each data source. This model is tuned using 11 hyperparameters. The third model combines CNN+CB by feeding the third-to-last layer of the CNN into the CB model, adding 784 features. For training, data is partitioned into 80% train and 20% test sets, stratified into 10 equal-width bins of wealth. A 4-fold cross-validation on the train set, stratified by SES (ses CV), is used to tune hyperparameters via Random Search on 200 combinations. This process is repeated three times independently with different random seeds, to control for fluctuations, and mean performance is reported. The models are evaluated on various configurations of training data, including different years of ground-truth data, relocation strategies, data augmentation ($*_{a}$), sample weight strategies ($*_{w}$), and feature sources, ensuring robust and comprehensive performance evaluation. For our analysis, we chose the model with the best performance (lowest root-mean-square error, RMSE) in each country: CB for Sierra Leone and Liberia, CBw for Uganda and Rwanda, and CNNa+CB for Gabon and South Africa.
M3: RWI inference model by Chi et al. 2022 [2, 5] utilizes $112$ features in total, including high-resolution satellite imagery, mobile phone network data, topographic maps, and aggregated, de-identified connectivity data from Facebook. These diverse data sources are processed through deep learning and computational algorithms to generate quantitative features for each village. These features are then used to train a gradient-boosted regression tree, a supervised machine-learning model, to predict RWI scores. Hyperparameters for the gradient-boosted tree are tuned to minimize the cross-validated MSE, using a grid search. They use three cross-validation approaches, two of which are also employed by M1: 5-fold cross-validation (basic CV), leave-one-country-out cross-validation (leave-country-out), and spatially stratified cross-validation (spatial CV). The use of different cross-validation methods helped highlight the potential upward bias in performance due to spatial auto-correlation in the training and test data. All analysis in this study is based on spatial CV, ensuring conservative and appropriate model training for geographic data with spatial auto-correlation.
@@
In summary, all models use a multi-modal and hybrid approach, combining features from various data sources and two model architectures, see Table 1 (bottom). M2 and M3 use satellite images by pre-training a CNN model and feeding the image embeddings into a gradient boosting algorithm. M1 also uses a hybrid approach but first pre-trains a gradient boosting algorithm to train a CNN, and then creates a feedback loop that enhances the initial model. Only M3 estimates the expected error of each location, though its explanatory power is limited (average $R^{2}=0.09$). All models predict the mean wealth of locations, but only M2 additionally predicts their standard deviation, and proposed two additional models to separate the effects of images and metadata.
@@ 2.3 Expected Trends
Before digging into the comparison of predictions between the three models of interest [2, 17, 8], it is crucial to acknowledge the discrepancies in the ground-truth data used to train them. Table 1 (top) outlines the specific years for which each approach utilized data for each country. This variation in time is significant because wealth and poverty are dynamic concepts that can fluctuate over time. To address this, rather than comparing directly the predicted values, we focus on the trends in predictions. We will achieve this by examining the trends of changes in Gross Domestic Product (GDP) per capita provided by the World Bank [34], as depicted in Figure 1 (top). By analyzing these trends, we can hypothesize that the comparison between the models will mirror the trends observed in GDP. This approach allows us to account for potential discrepancies arising from the distinct time-frames of the datasets.
@@
For our analysis, we perform pairwise comparisons between models, independently comparing M3 with both M1 and M2. When comparing M3 with M1, we apply a linear transformation to the RWI scores, converting them to their IWI equivalents in M1 due to their strong correlation (Pearson’s correlation coefficients: $0.93-0.99$ [17]). Similarly, we perform the corresponding transformation using M2 as the baseline. It is important to note that these transformations are approximate, as the RWI is inherently a relative score. Nevertheless, given the high-correlation between RWI and IWI, this assumption is justified. Future research could explore re-training M3 to directly predict IWI scores or transforming RWI into CWI, although this method remains imprecise [25]. The results comparing M3 with M1 and M2, presented in Section 0.A.3, should therefore be interpreted with caution.
Table 2: Prediction statistics. We show the number $N$ of target locations contained in the inferred poverty maps, and the mean, standard deviation (SD), Gini coefficient, and skewness test (SK) [32] of the predicted wealth scores for each model. None of the distributions are normal (SK test, $p<0.001$). Gini for M3 is omitted due to negative RWI values.
Country’s M1 (IWI) M2 (IWI) M3 (RWI)
Code $N$ Mean SD Gini SK $N$ Mean SD Gini SK $N$ Mean SD SK
SLE 13040 18.93 10.34 23.01 72.15 9881 21.87 6.09 11.54 70.52 8435 -0.35 0.29 44.13
LBR 16525 14.57 8.79 25.93 90.74 15597 23.35 3.74 7.75 75.15 6189 -0.38 0.33 36.12
UGA 27482 22.27 11.26 25.09 87.23 27791 41.83 5.21 5.95 86.16 25404 -0.23 0.31 65.00
RWA 23323 24.40 14.51 28.96 75.31 1150 47.79 3.46 3.74 13.65 3715 -0.04 0.36 24.85
ZAF 22687 61.14 18.10 16.85 4.12 3043 52.65 14.11 15.04 12.40 39072 -0.06 0.48 57.46
GAB 4835 21.89 8.86 17.93 46.60 1135 48.32 6.72 7.38 12.61 1110 -0.24 0.44 12.13
Table 3: Expected vs. observed wealth differences. Significant differences in predicted wealth between models are observed in LBR, UGA, RWA, and GAB, with M2 consistently predicting wealthier regions. RWA followed the expected trend but was notably over-represented. Identical wealth and Gini values were expected for UGA and ZAF due to the shared years of ground-truth data. However, M1 under-estimated wealth in UGA and slightly overestimated it in ZAF. In both cases, M1 generated more heterogeneous distributions compared to M2. All differences are statistically significant ($p<0.001$, Kolmogorov-Smirnov test [11]). Highlighted values show the correct trend (+/-). No available data (NA) for GAB (trained on data from other countries/years [17]) and for SLE, LBR, and RWA (due to the absence of official data).
Country’s Code M1 vs. M2
@@
SLE 0.00 -0.07 NA 11.47 0.62 ***
LBR 0.01 -0.23 NA 18.19 0.83 ***
UGA 0.00 -0.31 0 19.15 0.87 ***
RWA -0.05 -0.32 NA 25.22 0.85 ***
ZAF 0.00 0.07 0 1.81 0.30 ***
GAB NA -0.38 NA 10.55 0.95 ***
Expected IWI Difference $=(GDP_{Y_{M1}}-GDP_{Y_{M2}})/(GDP_{Y_{M1}}+GDP_{Y_{M2}})$
Observed IWI Difference $=(\mu_{IWI_{M1}}-\mu_{IWI_{M2}})/(\mu_{IWI_{M1}}+\mu_{IWI_{M2}})$
Expected Gini Difference $=(Gini_{Y_{M1}}-Gini_{Y_{M2}})$
Observed Gini Difference $=(Gini_{M1}-Gini_{M2})$
$Y_{M_{*}}$ denotes the year of the ground-truth data used to train model $M_{*}$, which is then used to select the corresponding indicator, GDP per capita or Gini coefficient, from the World Bank.